Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

aNNA and mark

Expert replies
by figs » Tue Apr 07, 2009 10:20 am
Anna and Mark are included in a group of 8 people. If 4 people are selected at random and without replacement from this group, what is the probability that Mark and not Anna will be selected?

A. 1/7
B. 2/7
C. 3/7
D. 4/7
E. 5/7


OA:B
Join the discussion
Source: — Problem Solving |

by scoobydooby » Tue Apr 07, 2009 10:52 am
if out of 4 people selected Mark is already selected (probability: 1) we need to select the rest 3 people from the remaining 6 people (8-mark-anna)


so 1*6C3/8C4=2/7

hence, B
Join the discussion

by figs » Wed Apr 08, 2009 3:05 am
can someone explain this exercise with another way. I'm interesting in know how to do with conditional probability.
Join the discussion

by dtweah » Sat Apr 11, 2009 1:36 pm
figs wrote:can someone explain this exercise with another way. I'm interesting in know how to do with conditional probability.
Let E be the event that the first 3 persons selected are neither Mark nor Ann

then M be the event of selecting M

A be the event of selecting A. Then A^c be the event she is not selected = 1-A

Probability of selecting 4 people = Number of ways of selecting first 3 x number of ways of selecting Mark given that Anna is not selected.
.
P(E)= 6C3/8/C4

P(M|A^c) = means that of the remaining 5 persons to choose from, Anna cannot be chosen and so cannot the others b/c we must have Mark.

Let's evaluate P(M|A^c) which you request.
P(M|A^c) = P(M intercept A^c) /P(A^c)

Given the indepdence of the events, this becomes
P(M|A^c) = P(M intercept A^c) /P(A^c) = P(M) P(A^c)/P(A^c) =P(M)

Now we know P(M) is a certainty since we already have the other three people and the probabability of certainty is 1.

That was the solution he earlier provided. I am just giving it the explanation you wanted.
Join the discussion

by alexhard » Sun Apr 12, 2009 1:40 pm
Simpler solution:

Probability mark will be selected: 1/2

Probability Anna WILL be selected = 6/7*5/6*4/5*3/4 = 3/7 (no need to actually do the calculation: 3 people out of 7 will be selected, so it's just directly 3/7)


Probability Anna WILL NOT be selected is therefore 1 - 3/7 = 4/7

Multiply the two: 1/2 * 4/7 = 2/7

There's no reason to dabble in combinatorics for this kind of stuff.
Join the discussion

by Jeff@TargetTestPrep » Mon Feb 12, 2018 4:42 pm
figs wrote:Anna and Mark are included in a group of 8 people. If 4 people are selected at random and without replacement from this group, what is the probability that Mark and not Anna will be selected?

A. 1/7
B. 2/7
C. 3/7
D. 4/7
E. 5/7
The calculation of the number of ways to select Mark but not Anna requires that we only consider the 6 remaining individuals because we don’t count Anna at all, and we already know that Mark has been selected. Thus, there are 6 available individuals to fill just 3 slots: 6C3 = 6!/[3!(6-3)!] = (6 x 5 x 4)/(3 x 2 x 1) = 20.

The number of ways to select 4 people from 8 is 8C4 = 8!/[4!(8-4)!] = (8 x 7 x 6 x 5)/(4 x 3 x 2 x 1) = 7 x 2 x 5 = 70.

So the probability that Mark will be selected and Anna will not be selected is 20/70 = 2/7.

Answer: B

Jeffrey Miller
Head of GMAT Instruction
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews
Join the discussion